OpenStudy (anonymous):

Is this right? Cylinder Volume Formula - Pie x r^2 x h

OpenStudy (paki):

yeah

OpenStudy (anonymous):

Yay! :) So... If I'm trying to find the height not the volume what do I do? :)

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

The radius of the sphere below is 8 cm. Which equation can be used to find V, the volume of the sphere? A. V = 3/4π(8^2) B. V = 3/4π(8^3) C. V = 4/3π(8^2) D. V = 4/3π(8^3) <---

OpenStudy (harsha19111999):

You can use the D option

OpenStudy (harsha19111999):

You there?